40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 920     Entry time: Tue Aug 10 19:22:19 2010
Author: Jenna 
Type: Laser 
Category: GYRO 
Subject: Cavity is locking again! 

Upon Rana's suggestion, Alastair adjusted the polarization of the beam going through the EOM using the half wave plate before the EOM and a polarizing beamsplitter to make sure it was exactly S-polarized and locked the half wave plate at the correct position. He also tightened up some of the SMA cables, some of which were a bit loose. As you can see there's no more DC-offset in the PDH error signal (blue).

Attachment 1: TEK00006.PNG  14 kB  | Hide | Hide all
TEK00006.PNG
ELOG V3.1.3-